# Co-Evolving Agents: Learning from Failures as Hard Negative

### v1

@@ Appendix A Datasets
WebShop (webshop) is a goal-oriented web navigation environment where an agent interacts with an e-commerce platform to fulfill natural language shopping instructions. A final reward in $[0,1]$ is assigned based on how well the purchased item matches the instruction in terms of attributes and price. For training, we use the expert and GPT-4 ReAct-style trajectories provided by eto, retaining only those with final rewards above $0.7$ to ensure quality.
@@
ScienceWorld (scienceworld) is a text-based embodied environment for conducting elementary science experiments across ten domains, including thermodynamics and electricity. Each task consists of multiple optional subgoals, and the final reward in $[0,1]$ reflects the proportion of subgoals completed. Following eto, we exclude the two longest task types and subsample test instances with more than ten variations. We evaluate generalization using the original test set of unseen variations (e.g., training on boiling water and testing on boiling lead) and use the development set for seen scenarios. Expert trajectories are drawn from the official dataset, with additional GPT-4–generated trajectories filtered by final rewards, as in eto.
@@
InterCodeSQL (intercode) is an interactive SQL querying environment where an agent must translate natural language questions into executable SQL programs through multi-turn interaction. The environment provides a binary reward, which we normalize to $[0,1]$, indicating whether the final query produces the correct result. We use the expert trajectories collected by IPR with GPT-4, retaining only those with reward $1.0$. Since the original dataset does not contain intermediate reasoning, each trajectory is augmented with ReAct-style (react) thought annotations generated by GPT-4, as in IPR.
@@ Appendix B Implementation Details
We adopt Llama-2-7B-Chat (llama2) as the primary base model for constructing LLM agents, following prior work (eto). All models are optimized with AdamW (adamw), and we use the same hyperparameter settings across datasets for fair comparison. During the supervised fine-tuning (SFT) stage, we train for 3 epochs with a batch size of 48 and a learning rate of 1e-5, using 3% linear warmup followed by a cosine decay schedule. The base agent then explores each training instance once to collect failure trajectories. In the subsequent optimization phase, we train with a batch size of 32 and a learning rate of 1e-6 to 5e-7, with the DPO scaling factor $\beta$ set to $0.1\text{-}0.5$. The number of optimization epochs is fixed to 3, and the number of co-evolution iterations is set to 3 for WebShop and ScienceWorld and 5 for InterCodeSQL. All experiments are conducted on 8 NVIDIA H100 GPUs with 80GB.
Appendix C Quality of the hard negatives
C.1 WebShop
@@ C.3 InterCodeSQL
Appendix D The effect of hard negatives on capturing task-relevant sub-skills

### v4

@@ Appendix A Datasets
WebShop (Yao et al., 2022) is a goal-oriented web navigation environment where an agent interacts with an e-commerce platform to fulfill natural language shopping instructions. A final reward in $[0,1]$ is assigned based on how well the purchased item matches the instruction in terms of attributes and price. For training, we use the expert and GPT-4 ReAct-style trajectories provided by Song et al. (2024), retaining only those with final rewards above $0.7$ to ensure quality.
@@
ScienceWorld (Wang et al., 2022) is a text-based embodied environment for conducting elementary science experiments across ten domains, including thermodynamics and electricity. Each task consists of multiple optional subgoals, and the final reward in $[0,1]$ reflects the proportion of subgoals completed. Following Song et al. (2024), we exclude the two longest task types and subsample test instances with more than ten variations. We evaluate generalization using the original test set of unseen variations (e.g., training on boiling water and testing on boiling lead) and use the development set for seen scenarios. Expert trajectories are drawn from the official dataset, with additional GPT-4–generated trajectories filtered by final rewards, as in Song et al. (2024).
@@
InterCodeSQL (Yang et al., 2023) is an interactive SQL querying environment where an agent must translate natural language questions into executable SQL programs through multi-turn interaction. The environment provides a binary reward, which we normalize to $[0,1]$, indicating whether the final query produces the correct result. We use the expert trajectories collected by Xiong et al. (2024) with GPT-4, retaining only those with reward $1.0$. Since the original dataset does not contain intermediate reasoning, each trajectory is augmented with ReAct-style (Yao et al., 2023) thought annotations generated by GPT-4, as in Xiong et al. (2024).
@@ Appendix B Implementation Details
We adopt Llama-2-7B-Chat (Touvron et al., 2023) as the primary base model for constructing LLM agents, following prior work (Song et al., 2024). All models are optimized with AdamW (Loshchilov and Hutter, 2017), and we use the same hyperparameter settings across datasets for fair comparison. During the supervised fine-tuning (SFT) stage, we train for 3 epochs with a batch size of 48 and a learning rate of 1e-5, using 3% linear warmup followed by a cosine decay schedule. The base agent then explores each training instance once to collect failure trajectories. In the subsequent optimization phase, we train with a batch size of 32 and a learning rate of 1e-6 to 5e-7, with the DPO scaling factor $\beta$ set to $0.1\text{-}0.5$. The number of optimization epochs is fixed to 3, and the number of co-evolution iterations is set to 3 for WebShop and ScienceWorld and 5 for InterCodeSQL. All experiments are conducted on 8 NVIDIA H100 GPUs with 80GB.
Appendix C Sample Weighting for Target Agent Training
To balance expert prediction supervision, we normalize the weights of all expert prediction pairs associated with a given instruction such that their total contribution sums to $1.0$. This prevents these pairs, originating from both the target agent and the failure agent, from disproportionately dominating the training signal while still encouraging alignment with high reward behaviors.
For failure failure comparisons, we disable the SFT term by setting $\lambda_{\text{SFT}}=0$ and rely solely on DPO updates. This avoids injecting incorrect supervised signals from suboptimal trajectories and ensures that these pairs retain full influence in shaping the preference gradients. Such weighting preserves the value of fine grained failure distinctions, which are crucial for refining the agent’s preference landscape.
Appendix D Quality of the Hard Negatives
D.1 WebShop
@@ D.3 InterCodeSQL
Appendix E The effect of hard negatives on capturing task-relevant sub-skills